Find the maximum value of the function: -4(x-3)(x+2)
Answer by stanbon(75887)

Find the maximum value of the function: -4(x-3)(x+2)
-----
y = -4(x^2-x-6)
y = -4x^2+4x+24
----
Max occurs when x = -b/(2a) = -4(2*-4)= 1/2
----
The maximum value = f(1/2) = -4(1/4)+4(1/2)+24 = -1+2+24 = 25
===============================================================
Cheers,
Stan H.